# Can we spot a fake?

### v1

@@ Abstract.
The problem of detecting fake data inspires the following seemingly simple mathematical question. Sample a data point $X$ from the standard normal distribution in $\mathbb{R}^{n}$. An adversary observes $X$ and corrupts it by adding a vector $rt$, where they can choose any vector $t$ from a fixed set $T$ of adversary’s “tricks”, and where $r>0$ is a fixed radius. The adversary’s choice of $t=t(X)$ may depend on the true data $X$. The adversary wants to hide the corruption by making the fake data $X+rt$ statistically indistinguishable from the real data $X$. What is the largest radius $r=r(T)$ for which the adversary can create an undetectable fake? We show that for highly symmetric sets $T$, the detectability radius $r(T)$ approximately equals twice the scaled Gaussian width of $T$. The upper bound actually holds for arbitrary sets $T$ and generalizes to arbitrary, non-Gaussian distributions of real data $X$. The lower bound may fail for not highly symmetric $T$, but we conjecture that this problem can be solved by considering the focused version of the Gaussian width of $T$, which focuses on the most important directions of $T$.
@@ 1. Introduction
This general question can be nontrivial even in stunningly simple scenarios. Imagine we sample a data point $X$ from the standard normal distribution on $\mathbb{R}^{n}$. An adversary observes $X$ and corrupts it by adding a vector $rt$, where they can choose any vector $t=t(X)\in T$ from a fixed set $T\subset\mathbb{R}^{n}$ of adversary’s “tricks”, and where $r>0$ is a fixed radius. Allowing the adversary’s choice of $t=t(X)$ to depend on $X$ allows the adversary may to manipulate different data in different ways. The adversary wants to hide the corruption by making the fake data $X+rt$ statistically indistinguishable from the real data $X$. When can the adversary succeed?
@@ 1.2. Related work
A lot of effort has been made to compute the detectability radius for an outsider adversary whose choice of the vector $t\in T$ that is not allowed to depend on $X$. Instead of (1.2), in the “outsider adversary” scenario the tester is satisfied with the following weaker goal: find a subset $A\subset\mathbb{R}^{n}$ that satisfies
@@ 1.4. Example: the adversary adds any vector
On the other hand, let us take $r=2\sqrt{n}$ and simplify reality a little by pretending that the Euclidean norm of $X$ is exactly $\sqrt{n}$. Then the adversary can make an undetectable fake by adding the vector $-2X\in rT$ to $X$. Indeed, this effectively flips the vector $X$ about the origin, which does not change the distribution of $X$ at all.
@@ 1.7. What about general sets $T$?
This latter result can fail miserably for general sets $T$: the scaled Gaussian width can hugely overestimate the detectability radius. A simple example is where $T$ consists of all unit vectors in $\mathbb{R}^{n}$ whose first coordinate is either $1/2$ or $-1/2$. An elementary calculation carried out in Section 5.1 shows that $\bar{w}(T)\asymp\sqrt{n}$ while $r(T)\asymp O(1)$.
@@ 2. When is a fake detectable?
Indeed, if this were not true, there would exist vectors $x,y\in K$ and $t\in T$ satisfying $x/2=t-y/2$, or $x+y=2t$. Multiplying by $t$ on both sides would give
@@
$\mathbb{P}\left\{X\in\frac{r}{2}K\right\}$ =$=\mathbb{P}\left\{\sup_{t\in T}\left\langle X,\frac{t}{\lVert t\rVert_{2}^{2}}\right\rangle<\frac{r}{2}\right\}\quad\text{(by definition of $K$)}$
@@ 3. When is a fake undetectable?
Definition 3.1 of a highly symmetric set $T$ implies that the answer to the question “does a given point $x\in\mathbb{R}^{n}$ belong to $T$?” depends only on the support of $x$ and whether the norm of $x$ is sufficiently large. Let $S$ denote the family of subsets of $\{1,\ldots,n\}$ that are supports of points in $T$, that is
@@
By definition of $T$, this implies that $-2X_{I}\in rT$, so there exists $t=t(X)\in T$ such that $-2X_{I}=rt$. Whenever the event (3.3) does not hold, set $I(X)=I_{0}$ and $t(X)=t_{0}$ for some arbitrary but fixed $I_{0}\in S$ and $t_{0}\in T$. Summarizing, we have constructed a random set $I=I(X)\in S$ and a random point $t=t(X)\in T$ such that
@@
Adding the vector $-2X_{I}$ to the vector $X$ is equivalent to reversing the signs of the coefficients of $X$ indexed by $I$. And since the choice of $I$ is determined by the absolute values of the coefficients of $X$ and is independent of their signs, reversing the coordinates of $X$ indexed by $I$ does not change the distribution of $X$. (This is a consequence of the symmetry of the Gaussian distribution, which we explain in detail in Lemma 3.3 below.) Hence $X-2X_{I}$ has the same distribution as $X$. Therefore, for any set $A\subset\mathbb{R}^{n}$ we have
$\mathbb{P}\{X\in A\}$ =$=\mathbb{P}\{X-2X_{I}\in A\}$
$\leq\mathbb{P}\{X-2X_{I}\in A\text{ and }-2X_{I}=rt\}+\mathbb{P}\{-2X_{I}\neq rt\}$
@@
Let $X=(X_{1},\ldots,X_{n})$ be a random vector in $\mathbb{R}^{n}$ whose coordinates $X_{i}$ are independent and have symmetric distributions. [^1] A random variable $\xi$ has symmetric distribution if $\xi$ has the same distribution as $-\xi$. Let $\theta=(\theta_{1},\ldots,\theta_{n})\in\{-1,1\}^{n}$ be a random vector whose value is determined by the absolute values of the coefficients of $X$. Then the random vector $(\theta_{1}X_{1},\ldots,\theta_{n}X_{n})$ has the same distribution as $X$.
@@ 4. Extensions for non-Gaussian data
Let $X$ be any random vector taking values in $\mathbb{R}^{n}$. Let $T\subset\mathbb{R}^{n}$ be any origin-symmetric set [^2] The assumption that $T$ is origin symmetric, i.e. that $T=-T$, is a bit annoying, but it is unavoidable in this statement, since a set $T$ consisting of a single point must satisfy $\bar{w}(T)=0$. However, it is easy to get around this assumption by replacing any set $T$ with the origin-symmetric set $T\cup-T$. and $u>0$ be any number. Assume that
@@
Let $X=(X_{1},\ldots,X_{n})$ be a random vector whose coordinates $X_{i}$ are independent, symmetric random variables taking values in the interval $[-1,1]$. Let $T\subset\mathbb{R}^{n}$ be a highly symmetric set and $u>0$ be any number. Assume that
@@ 5.1. The upper bound is not always sharp
If a set $T$ does not have enough symmetries, Theorem 3.2 can fail. Consider, for example, the set of all unit vectors whose first coordinate equals ±$\pm 1/2$, that is
@@ 5.2. Focused width
More generally, if $X$ is an arbitrary random vector taking values in $\mathbb{R}^{n}$, the focused $X$-width is defined as
@@ 5.3. Does focus help?
Indeed, if this were not true, there would exist vectors $x,y\in K$ and $t\in T$ such that $x/2=t-y/2$, or $x+y=2t$. By (5.4), we can find a vector $h\in H$ satisfying
@@
Multiplying by $h$ on both sides of the identity $x+y=2t$ would give
@@
$\mathbb{P}\left\{X\in\frac{r}{2}K\right\}$ =$=\mathbb{P}\left\{\sup_{h\in H}\left\langle X,\frac{h}{\lVert h\rVert_{2}^{2}}\right\rangle<\frac{r}{2}\right\}\quad\text{(by definition of $K$)}$

### v2

@@ Abstract.
The problem of detecting fake data inspires the following seemingly simple mathematical question. Sample a data point $X$ from the standard normal distribution in $\mathbb{R}^{n}$. An adversary observes $X$ and corrupts it by adding a vector $rt$, where they can choose any vector $t$ from a fixed set $T$ of the adversary’s “tricks”, and where $r>0$ is a fixed radius. The adversary’s choice of $t=t(X)$ may depend on the true data $X$. The adversary wants to hide the corruption by making the fake data $X+rt$ statistically indistinguishable from the real data $X$. What is the largest radius $r=r(T)$ for which the adversary can create an undetectable fake? We show that for highly symmetric sets $T$, the detectability radius $r(T)$ is approximately twice the scaled Gaussian width of $T$. The upper bound actually holds for arbitrary sets $T$ and generalizes to arbitrary, non-Gaussian distributions of real data $X$. The lower bound may fail for not highly symmetric $T$, but we conjecture that this problem can be solved by considering the focused version of the Gaussian width of $T$, which focuses on the most important directions of $T$.
@@ 1. Introduction
This general question can be nontrivial even in stunningly simple scenarios. Imagine we sample a data point $X$ from the standard normal distribution on $\mathbb{R}^{n}$. An adversary observes $X$ and corrupts it by adding a vector $rt$, where they can choose any vector $t=t(X)\in T$ from a fixed set $T\subset\mathbb{R}^{n}$ of adversary’s “tricks”, and where $r>0$ is a fixed radius. Allowing the adversary’s choice of $t=t(X)$ to depend on $X$ allows the adversary to manipulate different data in different ways. The adversary wants to hide the corruption by making the fake data $X+rt$ statistically indistinguishable from the real data $X$. When can the adversary succeed?
@@ 1.2. Related work
A lot of effort has been made to compute the detectability radius for an outsider adversary whose choice of the vector $t\in T$ is not allowed to depend on $X$. Instead of (1.2), in the “outsider adversary” scenario the tester is satisfied with the following weaker goal: find a subset $A\subset\mathbb{R}^{n}$ that satisfies
@@ 1.4. Example: the adversary adds any vector
On the other hand, let us take $r=2\sqrt{n}$ and simplify reality a little by pretending that the Euclidean norm of $X$ is exactly $\sqrt{n}$. Then the adversary can make an undetectable fake by adding the vector $2X\in rT$ to $X$. Indeed, this effectively flips the vector $X$ about the origin, which does not change the distribution of $X$ at all.
@@ 1.7. What about general sets $T$?
This latter result can fail miserably for general sets $T$: the scaled Gaussian width can hugely overestimate the detectability radius. A simple example is where $T$ consists of all unit vectors in $\mathbb{R}^{n}$ whose first coordinate is either $1/2$ or 1/2-1/2. An elementary calculation carried out in Section 5.1 shows that $\bar{w}(T)\asymp\sqrt{n}$ while $r(T)\asymp O(1)$.
@@ 2. When is a fake detectable?
Indeed, if this were not true, there would exist vectors $x,y\in K$ and $t\in T$ satisfying x/2=ty/2x/2=t-y/2, or $x+y=2t$. Taking the scalar product with $t$ on both sides would give
@@
$\mathbb{P}\left\{X\in\frac{r}{2}K\right\}$ $=\mathbb{P}\left\{\sup_{t\in T}\left\langle X,\frac{t}{\lVert t\rVert_{2}^{2}}\right\rangle<\frac{r}{2}\right\}\quad\text{(by definition of $K$)}$
@@ 3. When is a fake undetectable?
Definition 3.1 of a highly symmetric set $T$ implies that the answer to the question “does a given point $x\in\mathbb{R}^{n}$ belong to TT?” depends only on the support of $x$ and whether the norm of $x$ is sufficiently large. Let $S$ denote the family of subsets of $\{1,\ldots,n\}$ that are supports of points in $T$, that is
@@
By definition of $T$, this implies that $2X_{I}\in rT$, so there exists $t=t(X)\in T$ such that $2X_{I}=rt$. Whenever the event (3.3) does not hold, set $I(X)=I_{0}$ and $t(X)=t_{0}$ for some arbitrary but fixed $I_{0}\in S$ and $t_{0}\in T$. Summarizing, we have constructed a random set $I=I(X)\in S$ and a random point $t=t(X)\in T$ such that
@@
Adding the vector $2X_{I}$ to the vector $X$ is equivalent to reversing the signs of the coefficients of $X$ indexed by $I$. And since the choice of $I$ is determined by the absolute values of the coefficients of $X$ and is independent of their signs, reversing the coordinates of $X$ indexed by $I$ does not change the distribution of $X$. (This is a consequence of the symmetry of the Gaussian distribution, which we explain in detail in Lemma 3.3 below.) Hence $X-2X_{I}$ has the same distribution as $X$. Therefore, for any set $A\subset\mathbb{R}^{n}$ we have
$\mathbb{P}\{X\in A\}$ $=\mathbb{P}\{X-2X_{I}\in A\}$
[…]
@@
Let $X=(X_{1},\ldots,X_{n})$ be a random vector in $\mathbb{R}^{n}$ whose coordinates $X_{i}$ are independent and have symmetric distributions. [^1] A random variable $\xi$ has symmetric distribution if $\xi$ has the same distribution as $\xi$. Let $\theta=(\theta_{1},\ldots,\theta_{n})\in\{-1,1\}^{n}$ be a random vector whose value is determined by the absolute values of the coefficients of $X$. Then the random vector $(\theta_{1}X_{1},\ldots,\theta_{n}X_{n})$ has the same distribution as $X$.
@@ 4. Extensions for non-Gaussian data
Let $X$ be any random vector taking values in $\mathbb{R}^{n}$. Let $T\subset\mathbb{R}^{n}$ be any origin-symmetric set [^2] The assumption that $T$ is origin symmetric, i.e. that T=TT=-T, is a bit annoying, but it is unavoidable in this statement, since a set $T$ consisting of a single point must satisfy $\bar{w}(T)=0$. However, it is easy to get around this assumption by replacing any set $T$ with the origin-symmetric set $T\cup-T$. and $u>0$ be any number. Assume that
@@
Let $X=(X_{1},\ldots,X_{n})$ be a random vector whose coordinates $X_{i}$ are independent, symmetric random variables taking values in the interval [1,1][-1,1]. Let $T\subset\mathbb{R}^{n}$ be a highly symmetric set and $u>0$ be any number. Assume that
@@ 5.1. The upper bound is not always sharp
If a set $T$ does not have enough symmetries, Theorem 3.2 can fail. Consider, for example, the set of all unit vectors whose first coordinate equals $\pm 1/2$, that is
@@ 5.2. Focused width
To make it more broadly applicable, if $X$ is an arbitrary random vector taking values in $\mathbb{R}^{n}$, the focused $X$-width is defined as
@@ 5.3. Does focus help?
Indeed, if this were not true, there would exist vectors $x,y\in K$ and $t\in T$ such that x/2=ty/2x/2=t-y/2, or $x+y=2t$. By (5.4), we can find a vector $h\in H$ satisfying
@@
Taking the scalar product with $h$ on both sides of the identity $x+y=2t$ would give
@@
$\mathbb{P}\left\{X\in\frac{r}{2}K\right\}$ $=\mathbb{P}\left\{\sup_{h\in H}\left\langle X,\frac{h}{\lVert h\rVert_{2}^{2}}\right\rangle<\frac{r}{2}\right\}\quad\text{(by definition of $K$)}$
@@ 5.5. A conjecture
Finally, we conjecture that if detection is feasible for a particular value $r_{1}$, then it remains feasible for all values $r>r_{1}$; however, we currently lack a formal proof of this claim.
5.6. Acknowledgment
The authors sincerely thank the anonymous referees for their valuable comments, which helped improve this paper.